# Ren Xin1, Sheng Wang1, Yingbing Chen1, Jie Cheng2 and Ming Liu1,2

### v1

@@ III-G Training and Inference Details
A two-stage supervised learning process is conducted. We first train the prediction model from all the vehicles, like the general multi-agent training process. Then the map parameter and cost function training process is conducted by imitating the behavior demo of the ego vehicle. The raw context information is processed the same as in the first stage, and prediction results are inputs of the second stage.
@@ IV-A Implementation details
1) Dataset: Both the two training stages of our model are conducted on the Lyft Motion Prediction Dataset [34], which contains 134k, 25s time horizon, real-world driving samples, collected on a complex urban route in Palo Alto, California. HD maps and annotations for traffic signals/participants are provided at 10 Hz. A subset of about 3.6M frames is randomly selected for training and tested on a subset of the original validation set which contains 3600 frames.
2) Training: We train the prediction model using AdamW [35] optimizer using a batch size of 128 and a learning rate of 1e-4 for 18 epochs on a single RTX 3080 Lap-Top GPU. The planning RiskMap model is trained on 1/10 of the subset for one epoch on the same device. No dataset augmentation is implemented on either of the training stages. We train both the prediction model and RiskMap model to show trajectories for 3 seconds in the future, given the last 1.5 seconds of observation.
3) Evaluation: We run our model and other baselines on the logged data and report several common metrics including Average Distance Error(ADE), Final Distance Error(FDE) (Lat. and Lon. are short for lateral and longitudinal offset), Collision Rate, and Jerk$(m/s^{3})$. The average operation time of generating a planning result is 0.048s(Min. 0.017, Max. 0.111). The operating time meets the real-time requirement.
@@ IV-B Comparison with Baselines
We compare the planning results against two representative baselines on the Lyft dataset. ChauffeurNet[36], where the input is rasterized images and perturbation is used to deal with covariate shift. ChaufferNet-Vec, where the rasterized image input is replaced with the recent more advanced vectorized representation [10]. The comparison result are listed in Table I. The baselines are selected to show imitation learning performance. From where we know our method has the lowest Jerk value as it is strictly following dynamic rules and the sampling module offers Jerk optimized trajectories.
@@ IV-C Ablation Study of Model Components
Table II compares the performance of different module features. Risk is our proposed method of mapping Euler distance to risk space by (4). LN is to replace the Probabilistic module with a Linear scalar and bias. TV is to use temporal parameters on the planning time horizon. No check mark under this item means the model adopts a constant parameter along each time step of sampled trajectories. Different sampling count is tested to verify the extendability of our model. The sampling count only changes the sampling density of a fixed speed and steering range. From rows 1, 2, and 4(our standard model), we can see our proposed method with the same sampling count performs the best on ADE, FDE(Lon.), and Jerk metrics. The time variational mapping params can increase the performance of AED by 7.4%. And the ADE incrementation of the probabilistic model is 84.2% w.r.t Linear model. Rows 3,4,5 shows the performance of the map model will be influenced by downstream planning methods, higher sampling density will achieve better performance without additional training process, but users need to balance the performance and calculation resources.
@@ IV-D Ablation Study of Loss Items
The results of the Loss items ablation study are listed in Table III. Only one item is removed in each ablation study. The baseline model which adopts all the loss items is listed in row 4 of Table II. Removing Ind. Vel. Reg. means the model change the mode of learning target from ground truth to directly learn from sampled costs, as velocity difference is also a cost item. The result shows training loss gets divergent. From the table we find Cross Entropy loss can improve ADE performance by 22.2%, Cost Consistency loss item makes the progress of 14.8%. Cost Regression, in theory, will make the results of the model more robust, but in our small data quantity experiments, we obtain better results than the baseline model of 4.6%. Minimizing Demo cost is an intuitive loss item but it actually draws back the ADE performance of 9.2%. This target may lead to a local minimum of the optimization process of the model.
@@ V Conclusion
In this work, we found an efficient multi-modal prediction result utilization method for local planning by unifying it together with other driving contexts and diagnosing it as RiskMap. An interpretable and hands-free parameter tuning planner is adopted to demonstrate the utilization and efficiency of RiskMap. We proposed the training and planning framework and analyzed the efficiency of model modules and training tools. This model can be further used in driver behavior analysis and customized agent behavior in simulators. However, the effectiveness of the use of optimization-based methods needs to be validated in future works. The learning-based planners are also supposed to have better generalizability as RiskMap offers extracted expressions of map information.
Method (5s) Planning Error(m) Prediction Error(m)
ADE FDE ADE FDE
MM-Decoder 1.6737 4.9665 1.3968 3.3180
DIPP 1.6945 4.2801 1.4791 3.4029
Vector Field(ours) 1.6642 4.3923 1.3830 3.1595

### v3

@@ III-G Training and Inference Details
A two-stage supervised learning process is conducted. We first train the prediction model from all the vehicles, like the general multiagent training process. Then the map parameter and cost function training process is conducted by imitating the behavior demo of the ego vehicle. The raw context information is processed the same as in the first stage, and prediction results are input into the second stage.
@@ IV-A Implementation Details
(1) Dataset: Both the two training stages of our model are carried out on the Lyft Motion Prediction Dataset [29], which contains 134 k, 25 s time horizon, real-world driving samples, collected on a complex urban route in Palo Alto, California. HD maps and annotations for traffic signals/participants are provided at 10 Hz. A subset of about 3.6 M frames is randomly selected for training and tested on a subset of the original validation set which contains 3600 frames.
(2) Training: We train the prediction model using the AdamW [30] optimizer using a batch size of 128 and a learning rate of 0.0001 for 18 epochs on a single RTX 3080 laptop GPU. The RiskMap planning model is trained on 1/10 of the subset for one epoch on the same device. No dataset augmentation is implemented in either of the training stages. We train both the prediction model and RiskMap model to show trajectories for 3 s in the future, given the last 1.5 s of observation.
(3) Evaluation: We run our model and other baselines on the logged data and report several common metrics including Average Distance Error (ADE), Final Distance Error (FDE), Lat. and Lon. are short for lateral and longitudinal offset, Collision Rate, and Jerk. The average operation time of generating a planning result is 0.048 s (Min. 0.017, Max. 0.111). The operating time meets the real-time requirement.
@@ IV-B Comparison with Baselines
We compare the planning results against two representative baselines on the Lyft dataset. ChauffeurNet [31], where the input is rasterized images and perturbation is used to deal with co-variate shift. ChaufferNet-Vec, where the rasterized image input is replaced with the more advanced vectorized representation [11]. The comparison result are listed in Table I. The baselines are selected to show imitation learning performance. From Table I we know our method has the lowest jerk value, as it is strictly following dynamic rules and the sampling module offers jerk optimized trajectories. The capability of dealing with collision and copying expert behavior is not surpassing learning-based methods no matter what kind of cost remapping method is adopted. The results illustrate that pure distance information may not sufficient for drivers decision. But it still achieves comparable performance with better interpretability.
@@ IV-C Ablation Study of Model Components
Table II compares the performance of different module features. Risk is our proposed method of mapping Euler distance to risk space by (4). LN is to replace the Probabilistic module with a Linear scalar and bias. TV is to use temporal parameters on the planning time horizon. No check mark under this item means the model adopts a constant parameter along each time step of sampled trajectories. Different sampling count is tested to verify the expendability of our model. The sampling count only changes the sampling density of a fixed speed and steering range. From rows 1, 2, and 4, we can see that our proposed method with the same sampling count performs the best on ADE, FDE (Lon.), and jerk metrics. The time variational mapping parameters can increase the performance of AED by 7.4%. And the ADE incrementation of the probabilistic model is 84.2% w.r.t. linear model. Rows 3, 4, and 5 show that the performance of the map model will be influenced by downstream planning methods, higher sampling density will achieve better performance without additional training process, but users need to balance the performance and computation resources.
@@ IV-D Ablation Study of Loss Items
The results of the Loss items ablation study are listed in Table III. Only one item is removed in each ablation study. The baseline model which adopts all the loss items is listed in row 4 of Table II. Removing Individual Velocity Regression means that the model removes the loss of directly learning velocity value from ground truth as introduced in (7). The results shows that our model convergent in freezing robot status. From the table we find Cross Entropy loss can improve ADE performance by 22.2%, Cost Consistency loss item makes the progress of 14.8%. Cost Regression, in theory, will make the results of the model more robust, but in our small data quantity experiments, we obtain better results than the baseline model of 4.6%. Minimizing Demo Cost is an intuitive loss item, but the ADE performance increases by 9.2% with the absence of it. This target may lead to a local minimum in the optimization process of the model.
@@ V Conclusion
In this work, we found an efficient multi-modal prediction result utilization method for local planning by unifying it together with other driving contexts and diagnosing it as RiskMap. An interpretable and hands-free parameter tuning planner is adopted to demonstrate the utilization and efficiency of RiskMap. We propose the training and planning framework and analyze the efficiency of model modules and training tools. This model can be further used in driver behavior analysis and customized agent behavior in simulators. However, the effectiveness of the use of cost-map-based methods needs to be validated in future works. The learning-based planners are also supposed to have better generalizability as RiskMap offers extracted expressions of map information.